# The Mean of Multi-Object Trajectories

### v1

@@ III-B1 Problem Transformation
If $y^{*}$ minimizes $g$, then $\mathcal{R}(y^{*})$ minimizes $f$, and $g(y^{*})=f(\mathcal{R}(y^{*}))$. (See SM Subsection I-B for proof)
Using this equivalent representation, the Fréchet mean trajectory (3) is given by the solution to Problem (8) below on the space $\mathbb{B}^{|\mathbb{K}|}\times\mathbb{X}^{|\mathbb{K}|}$. This is stated more concisely in the following (see SM Subsection I-C for proof).
@@
[…]
@@
[…]
@@
For any $\gamma\in\mathbb{B}^{|\mathbb{K}|}$, let $[x_{1}^{*}(\gamma),...,x_{|\mathbb{K}|}^{*}(\gamma)]\triangleq\arg\min_{x\in\mathbb{X}^{|\mathbb{K}|}}U^{(r)}(\gamma,x)$. If $\gamma_{k}\!=\!0$ or $\ \mathbf{1}_{\mathcal{D}_{\!v^{(n)}\!}}^{(k)}=\!0$, for all $n=\!1:N$, then $x_{k}^{*}(\gamma)\!=\!\boldsymbol{0}\in\mathbb{X}$. Further, if $\gamma_{k}\!=\!1$, then
$x_{k}^{*}(\gamma)=\arg\min_{x_{k}\in\mathbb{X}}{\displaystyle\sum_{n:D_{\!v^{(n)}\!}\ni k}}\frac{[d_{\mathbb{X}}^{(c)\!}(x_{k},v^{(n)\!}(k))]^{r}}{|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|^{r}}.$ (15)
This result (see SM Subsection I-D for proof) provides an explicit expression for each component of $x^{*}(\gamma)$. Specifically, if at time $k$, the state of $x^{*}(\gamma)$ does not exist ($\gamma_{k}\!=\!0$) or all states of the sample trajectories do not exist, then $x_{k}^{*}(\gamma)$ is set to $\boldsymbol{0}\in\mathbb{X}$. If the state of $x^{*}(\gamma)$ exists at time $k$ ($\gamma_{k}\!=\!1$), then $x_{k}^{*}(\gamma)$ is given by the $r^{th}$-order Fréchet mean of the sample trajectory states that exist, weighted by $|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|^{r}$. Note that $x_{k}^{*}(\gamma)$ in (15) may not be unique, but any minimizer of the cost would suffice.
@@ III-B2 Mean Trajectory via Greedy Search
Targeting efficiency to suit online applications, we present a greedy search algorithm that generates a new $\tilde{\gamma}$ from $\gamma$ simply by changing the component $\gamma_{k}\in\{0,1\}$ to its complement (i.e., from 0 to 1 or vice-versa) if this reduces the cost $W^{(r)}$. This process is cycled through all values of $k\in\mathbb{K}$ until the cost reduction falls below some threshold. The pseudocode is given in Algorithm 1. If the complexity of solving Problem (15) is $\mathcal{O}(M)$, then the worst case complexity of this algorithm is $\mathcal{O}(T_{t}K^{2}M)$. The first row of Figure 1 in Section I shows the mean trajectories computed using this method.
@@
$\gamma^{*}:=\tilde{\gamma}$; $\delta_{\textrm{cost}}\!:=\!$ $\textrm{cost}^{*}-W^{(r)}(\tilde{\gamma})$; $\textrm{cost}^{*}:=W^{(r)}(\tilde{\gamma})$;
@@
Observe from (9)-(13) that the dependence of trajectory $x^{*}(\gamma)$ on $\gamma$ is captured in the factors $\left(|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|^{r}\right)^{-1}$, $n\in\{1:N\}$. In practice, these factors have similar values for different $n$ since the estimates of the same trajectory, but by different nodes, would have similar lengths. Moreover, these lengths are relatively long, thereby diminishing the influence of $\left(|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|^{r}\right)^{-1}$ on $x^{*}(\gamma)$. Hence, to further reduce computations, the optimal $x^{*}(\gamma)$ can be approximated by an $x^{*}$ with components
@@ IV-A The Fréchet Mean Multi-Object Trajectory
Consider $r\geq 1$, $p\geq 0$, $c\in[p,\sqrt[r]{2}p]$, and the OSPA trajectory distance $d_{\mathbb{T}}^{(c)}$ on $\mathbb{T}$ given by [^3] We can use $d_{\mathbb{T}}^{(c)}(y,z)=\min\{d_{\mathbb{T}}^{(q)}(y,z),c\}$, where $d_{\mathbb{T}}^{(q)}(y,z)$ is the OSPA trajectory distance, but for convenience we set $q=c$ (2). For any two multi-object trajectories $\boldsymbol{X}=\{\boldsymbol{X}_{1:m}\}$ and $\boldsymbol{Y}=\{\boldsymbol{Y}_{1:s}\}$ (without loss of generality, assume $m\leq s$), the OSPA(2) (multi-object trajectory) distance of order $r$, cardinality penalty $p$ and cut-off $c$ is defined as [muller2020metrics, schuhmacher2008consistent, beard2020solution]
$\!\!\!\!\boldsymbol{d}(\boldsymbol{X}\!,\boldsymbol{Y})=\left[\min_{\pi\in\Pi_{s}}\!\sum_{\ell=1}^{m}\!\frac{[d_{\mathbb{T}}^{(c)}\!(\boldsymbol{X}_{\!\ell},\boldsymbol{Y}_{\!\!\pi(\ell)})]^{r}+|s\!-\!m|p^{r}}{s}\!\right]^{\!\frac{1}{r}}\!\!,$ (18)
with $\boldsymbol{d}(\emptyset\!,\emptyset)=0$, where $\Pi_{s}$ denotes the space of permutations on $\{1:s\}$. If $m>s$, we define $\boldsymbol{d}(\boldsymbol{X}\!,\boldsymbol{Y})=\boldsymbol{d}(\boldsymbol{Y}\!,\boldsymbol{X})$. Note that for simplicity, we suppressed the dependence on the parameters $r,$ $c$ and $p$.
@@
More generally $\boldsymbol{d}(\boldsymbol{X}\!,\boldsymbol{Y})$ is the OSPA distance between the multi-object trajectories $\boldsymbol{X}\!,\boldsymbol{Y}$ when $d_{\mathbb{T}}^{(c)}(y,z)=\min\{d_{\mathbb{T}}(y,z),c\}$, for any trajectory distance $d_{\mathbb{T}}(y,z)$. Note also that the OSPA construction admits variations of the OSPA metric such as the COLA metric [barrios2015cardinalized], and the TT metric [muller2020metrics] by extending (18) to
@@
An upper bound on the cardinality of the Fréchet mean multi-object trajectory w.r.t. the OSPA-based metrics is given in the following proposition. The proof is given in SM Subsection I-E.
@@ IV-B1 Problem Transformation
Since we are interested in computing the mean of the multi-object trajectories $\boldsymbol{Y}^{(1:N)}$, the list lenghth $L$ should not be less than the cardinality of this mean. Using the upper bound on the Fréchet mean in Proposition 4, we fix $L=\sum_{n=1}^{N}|\boldsymbol{Y}^{(n)}|$.
@@
Using the transformation $\mathcal{A}$ above, Problem (17)-(LABEL:eq:ospa-based-traj-ori) on $\mathcal{M}(\mathbb{T})$ is equivalent to Problem (23) on $\mathbb{B}^{L}\times\mathbb{T}^{L}\times\left[\Pi_{L}\right]^{N}$. This is stated more concisely in the following (see SM Subsection I-F for proof).
@@
$Q_{\omega}^{(r)}(\eta,\tau)$ $=\sum_{\ell=1}^{L}\eta_{\ell}\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})\text{+}(1-\eta_{\ell})\bar{\phi}_{\omega,\ell}^{(r)}(\eta),$ (20)
$\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})$ $=\sum_{n=1}^{N}\!\frac{\xi_{\omega(n,\ell)}^{(n)}[d_{\mathbb{T}}^{(c)\!}(\tau_{\ell},\chi_{\omega(n,\ell)}^{(n)})]^{r}+(1\!-\!\xi_{\omega(n,\ell)}^{(n)})p^{r}}{\kappa(|\!|\eta|\!|_{1},|\!|\xi^{(n)}|\!|_{1})},$ (21)
$\bar{\phi}_{\omega,\ell}^{(r)}(\eta)$ $=\sum_{n=1}^{N}\!\frac{\xi_{\omega(n,\ell)}^{(n)}p^{r}}{\kappa(|\!|\eta|\!|_{1},|\!|\xi^{(n)}|\!|_{1})},$ (22)
@@
$(\hat{\eta},\hat{\tau})$ $=\arg\min_{(\eta,\tau)\in\mathbb{B}^{L}\times\mathbb{T}^{L}}\min_{\omega\in\Omega^{(1:N)}(\eta)}Q_{\omega}^{(r)}(\eta,\tau),$ (23)
@@
For any $\eta\!\in\!\mathbb{B}^{L}$ and $\omega\!\in\!\Omega^{(1:N)}(\eta)$, let $[\tau_{1}^{*}(\eta,\omega),...,\tau_{L}^{*}(\eta,\omega)]=\arg\min_{\tau\in\mathbb{T}^{L}}Q_{\omega}^{(r)}(\eta,\tau)$. If $\eta_{\ell}=0$, or $\xi_{\omega(n,\ell)}^{(n)}=0$ for all $n=\!1\!:\!N$, then $\tau_{\ell}^{*}(\eta,\omega)\!=\!\boldsymbol{0}$. Otherwise,
$\tau_{\ell}^{*}(\eta,\omega)\!=\!\arg\min_{\tau_{\ell}\in\mathbb{T}}\!\sum_{n:\xi_{\omega(n,\ell)}^{(n)}=1}\!\!\frac{d_{\mathbb{T}}^{(c)}\!(\tau_{\ell},\chi_{\omega(n,\ell)}^{(n)})^{r}}{\kappa(|\!|\eta|\!|_{1},|\!|\xi^{(n)}|\!|_{1})}.$ (27)
This result (see SM Subsection I-G for proof) provides an explicit expression for each component trajectory $\tau_{\ell}^{*}(\eta,\omega)$ of the trajectory list $\tau^{*}(\eta,\omega)$ that minimizes $Q_{\omega}^{(r)}(\eta,\cdot)$. Specifically, the $\ell^{th}$ trajectory $\tau_{\ell}^{*}(\eta,\omega)$ is set to $\boldsymbol{0}\in\mathbb{T}$, if $\eta_{\ell}=0$ or $\tau_{\ell}^{*}(\eta,\omega)$ is not associated (via $\omega$) with any sample trajectories. Otherwise, $\tau_{\ell}^{*}(\eta,\omega)$ is the $r^{th}$-order Fréchet mean of the associated sample trajectories $\chi_{\omega(1,\ell)}^{(1)}$, …, $\chi_{\omega(N,\ell)}^{(N)}$ that exist, weighted by $\kappa(|\!|\eta|\!|_{1},|\!|\xi^{(1)}|\!|_{1})$, …, $\kappa(|\!|\eta|\!|_{1},|\!|\xi^{(N)}|\!|_{1})$, which can be computed using the method in Subsection III-B2. Note that the solution of Problem (27) might not be unique. Nonetheless, any minimizer of the cost would be sufficient.
@@ IV-B2 Mean Multi-Object Trajectory via Greedy Search
[…]
@@
The minimization on $\mathbb{B}^{L}$ is carried out by optimizing each coordinate of $\eta$ through the functions $\mathtt{addTrajectories}$ (Algorithm 4) and $\mathtt{deleteTrajectories}$ (Algorithm 5). The $\mathtt{addTrajectories}$ function updates each coordinate of $\eta$ from 0 to 1 whenever such a change reduces the cost. The $\mathtt{deleteTrajectories}$ function updates each coordinate of $\eta$ from 1 to 0 if doing so results in a cost reduction. If the complexity of the algorithm to solve Problem (27) is $\mathcal{O}(Z)$, the worst case complexity of Algorithm 2 is $\mathcal{O}(T_{m}L(NL^{2}+Z))$. In the second row of Figure 1 in Section I, we show examples of multi-object trajectory consensus of three point patterns computed by this algorithm.
@@
Find ´ such that $\textrm{Find }\acute{\ell}\textrm{ such that }\acute{\omega}(n,\acute{\ell})=i$;
$\acute{i}:=\acute{\omega}(n,\ell)$; $\acute{\omega}(n,\ell):=i$$\acute{\omega}(n,\acute{\ell}):=\acute{i}$;

### v2

@@ III-B1 Problem Transformation
If $y^{*}$ minimizes $g$, then $\mathcal{R}(y^{*})$ minimizes $f$, and $g(y^{*})=f(\mathcal{R}(y^{*}))$. (See SM Subsection I-C for proof)
Using this equivalent representation, the Fréchet mean trajectory (3) is given by the solution to Problem (8) below on the space $\mathbb{B}^{|\mathbb{K}|}\times\mathbb{X}^{|\mathbb{K}|}$. This is stated more concisely in the following (see SM Subsection I-D for proof).
@@
$U^{(r)}(\gamma,x)$ $\!\!\!\!=\!\!\!\!$ ${\displaystyle\sum_{k\in\mathbb{K}}}\gamma_{k}\psi_{k}^{(r)}(\gamma,x_{k})+\left(1-\gamma_{k}\right)\bar{\psi}_{k}^{(r)}(\gamma),$ (9)
$\!\!\!\psi_{k}^{(r)}(\gamma,x_{k})$ $\!\!\!\!=\!\!\!\!$ ${\displaystyle\sum_{n=1}^{N}\!}\frac{\mathbf{1}_{\mathcal{D}_{\!v^{(n)}\!}}^{(k)}[d_{\mathbb{X}}^{(c)\!}(x_{k},v^{(n)\!}(k))]^{r}\!+\!\mathbf{1}_{\mathcal{\overline{D}}_{\!v^{(n)}\!}}^{(k)}c^{r}}{|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|},$ (10)
$\bar{\psi}_{k}^{(r)}(\gamma)$ $\!\!\!\!=\!\!\!\!$ ${\displaystyle\sum_{n=1}^{N}}\frac{\mathbf{1}_{\mathcal{D}_{\!v^{(n)}}}^{(k)}c^{r}}{|\mathcal{D}_{\gamma}\cup\mathcal{D}_{\!v^{(n)}}|},$ (11)
$\mathcal{D}_{\gamma}$ $\!\!\!\!=\!\!\!\!$ $\{k\in\mathbb{K}:\gamma_{k}=1\}.$ (12)
@@
$x^{*}(\gamma)$ $\!\!\!\!\triangleq\!\!\!\!$ $\arg\min_{x\in\mathbb{X}^{|\mathbb{K}|}}U^{(r)}(\gamma,x),$ (13)
$W^{(r)}(\gamma)$ $\!\!\!\!\triangleq\!\!\!\!$ $U^{(r)}(\gamma,x^{*}(\gamma)).$ (14)
@@
[…]
$x_{k}^{*}(\gamma)=\arg\min_{x_{k}\in\mathbb{X}}{\displaystyle\sum_{n:\mathcal{D}_{\!v^{(n)}\!}\ni k}}\frac{[d_{\mathbb{X}}^{(c)\!}(x_{k},v^{(n)\!}(k))]^{r}}{|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|}.$ (15)
This result (see SM Subsection I-E for proof) provides an explicit expression for each component of $x^{*}(\gamma)$. Specifically, if at time $k$, the state of $x^{*}(\gamma)$ does not exist ($\gamma_{k}\!=\!0$) or all states of the sample trajectories do not exist, then $x_{k}^{*}(\gamma)$ is set to $\boldsymbol{0}\in\mathbb{X}$. If the state of $x^{*}(\gamma)$ exists at time $k$ ($\gamma_{k}\!=\!1$), then $x_{k}^{*}(\gamma)$ is given by the $r^{th}$-order Fréchet mean of the sample trajectory states that exist, weighted by $|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|$. Note that $x_{k}^{*}(\gamma)$ in (15) may not be unique, but any minimizer of the cost would suffice.
@@ III-B2 Mean Trajectory via Greedy Search
Targeting efficiency to suit online applications, we present a greedy search algorithm that generates a new $\tilde{\gamma}$ from $\gamma$ simply by changing the component $\gamma_{k}\in\{0,1\}$ to its complement (i.e., from 0 to 1 or vice-versa) if this reduces the cost $W^{(r)}$. This process is cycled through all values of $k\in\mathbb{K}$ until the cost reduction falls below some threshold. The pseudocode is given in Algorithm 1. If the complexity of solving Problem (15) is $\mathcal{O}(M)$, then the worst case complexity of this algorithm is $\mathcal{O}(T_{t}K^{2}M)$. Figure 1 shows the mean trajectories computed using this method.
@@
if $\textrm{$W^{(r)}$($\tilde{\gamma}$)}<\textrm{cost}^{*}$ then
@@
Observe from (9)-(13) that the dependence of trajectory $x^{*}(\gamma)$ on $\gamma$ is captured in the factors $\left(|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|\right)^{-1}$, $n\in\{1:N\}$. In practice, these factors have similar values for different $n$ since the estimates of the same trajectory, but by different nodes, would have similar lengths. Moreover, these lengths are relatively long, thereby diminishing the influence of $\left(|\mathcal{D}_{\gamma}\cup\mathcal{D}_{v^{(n)}}|\right)^{-1}$ on $x^{*}(\gamma)$. Hence, to further reduce computations, the optimal $x^{*}(\gamma)$ can be approximated by an $x^{*}$ with components
@@ IV-A The Fréchet Mean Multi-Object Trajectory
Consider $r\geq 1$, $p\geq 0$, $c\in[p,\sqrt[r]{2}p]$, and the OSPA trajectory distance $d_{\mathbb{T}}^{(c,r)}$ on $\mathbb{T}$ given by [^5] We can use $d_{\mathbb{T}}^{(c,r)}(y,z)=\min\{d_{\mathbb{T}}^{(q,l)}(y,z),c\}$, where $d_{\mathbb{T}}^{(q,l)}(y,z)$ is the OSPA trajectory distance, but for convenience we set $q=c$ and $l=r$. (2). For any two multi-object trajectories $\boldsymbol{X}=\{\boldsymbol{X}_{1:m}\}$ and $\boldsymbol{Y}=\{\boldsymbol{Y}_{1:s}\}$ (without loss of generality, assume $m\leq s$), the OSPA(2) (multi-object trajectory) distance of order $r$, cardinality penalty $p$ and cut-off $c$ is defined as [muller2020metrics, schuhmacher2008consistent, beard2020solution]
$\boldsymbol{d}(\boldsymbol{X}\!,\boldsymbol{Y})=\left[\min_{\pi\in\Pi_{s}}\!\sum_{\ell=1}^{m}\!\frac{[d_{\mathbb{T}}^{(c,r)}\!(\boldsymbol{X}_{\!\ell},\boldsymbol{Y}_{\!\!\pi(\ell)})]^{r}+|s\!-\!m|p^{r}}{s}\!\right]^{\!\frac{1}{r}}\!\!,$ (18)
with $\boldsymbol{d}(\emptyset\!,\emptyset)=0$, where $\Pi_{s}$ denotes the space of permutations on $\{1:s\}$. If $m>s$, we define $\boldsymbol{d}(\boldsymbol{X}\!,\boldsymbol{Y})=\boldsymbol{d}(\boldsymbol{Y}\!,\boldsymbol{X})$. Note that for simplicity, we suppressed the dependence on the parameters $r,$ $c$ and $p$. This definition is also based on the fact that the OSPA metric [schuhmacher2008consistent] can be extended to the case $c\in[p,\sqrt[r]{2}p]$, termed the relative TT metric in [muller2020metrics] (see Section 2 of the referred paper for more details).
@@
More generally $\boldsymbol{d}(\boldsymbol{X}\!,\boldsymbol{Y})$ is the OSPA distance between the multi-object trajectories $\boldsymbol{X}\!,\boldsymbol{Y}$ when $d_{\mathbb{T}}^{(c)}(y,z)=\min\{d_{\mathbb{T}}(y,z),c\}$, for any trajectory distance $d_{\mathbb{T}}(y,z)$. Note also that the OSPA construction admits variations of the OSPA metric such as the COLA metric [barrios2017metrics], and the TT metric [muller2020metrics] by extending (18) to
@@
An upper bound on the cardinality of the Fréchet mean multi-object trajectory w.r.t. the OSPA-based metrics is given in the following proposition. The proof is given in SM Subsection I-F.
@@ IV-B1 Problem Transformation
Since we are interested in computing the mean of the multi-object trajectories $\boldsymbol{Y}^{(1:N)}$, the list length $L$ should not be less than the cardinality of this mean. Using the upper bound on the Fréchet mean in Proposition 4, we fix $L=\sum_{n=1}^{N}|\boldsymbol{Y}^{(n)}|$.
@@
Using the transformation $\mathcal{A}$ above, Problem (17)-(LABEL:eq:ospa-based-traj-ori) on $\mathcal{M}(\mathbb{T})$ is equivalent to Problem (23) on $\mathbb{B}^{L}\times\mathbb{T}^{L}\times\left[\Pi_{L}\right]^{N}$. This is stated more concisely in the following (see SM Subsection I-G for proof).
@@
$Q_{\omega}^{(r)}(\eta,\tau)$ $=\sum_{\ell=1}^{L}\eta_{\ell}\phi_{\omega,\ell}^{(r)}(\eta,\tau_{\ell})+(1-\eta_{\ell})\bar{\phi}_{\omega,\ell}^{(r)}(\eta),$ (20)
[…]
@@
[…]
@@
For any $\eta\!\in\!\mathbb{B}^{L}$ and $\omega\!\in\!\Omega^{(1:N)}(\eta)$, let $\tau_{1}^{*}(\eta,\omega),...,\tau_{L}^{*}(\eta,\omega)]=\arg\min_{\tau\in\mathbb{T}^{L}}Q_{\omega}^{(r)}(\eta,\tau)$. If $\eta_{\ell}=0$, or $\xi_{\omega(n,\ell)}^{(n)}=0$ for all n=1:Nn=\!1\!:\!N, then $\tau_{\ell}^{*}(\eta,\omega)\!=\!\boldsymbol{0}$. Otherwise,
[…]
@@ IV-B2 Mean Multi-Object Trajectory via Greedy Search
$\eta^{*}:=\eta$; $\omega^{*}$:= $\omega$; $\delta_{\textrm{cost}}:=$ $\textrm{cost}^{*}$- cost;  $\textrm{ cost}^{*}$:= cost;
@@
The minimization on $\mathbb{B}^{L}$ is carried out by optimizing each coordinate of $\eta$ through the functions $\mathtt{addTrajectories}$ (Algorithm 4) and $\mathtt{deleteTrajectories}$ (Algorithm 5). The $\mathtt{addTrajectories}$ function updates each coordinate of $\eta$ from 0 to 1 whenever such a change reduces the cost. The $\mathtt{deleteTrajectories}$ function updates each coordinate of $\eta$ from 1 to 0 if doing so results in a cost reduction. If the complexity of the algorithm to solve Problem (27) is $\mathcal{O}(Z)$, the worst case complexity of Algorithm 2 is $\mathcal{O}(T_{m}L(NL^{2}+Z))$. In Figure 2, we show examples of multi-object trajectory consensus of three point patterns computed by this algorithm.
@@
$\textrm{Find }\acute{\ell}\textrm{ such that }\acute{\omega}(n,\acute{\ell})=i$;
$\acute{i}:=\acute{\omega}(n,\ell)$; $\acute{\omega}(n,\ell):=i$; $\acute{\omega}(n,\acute{\ell}):=\acute{i}$;